[Brahms-l] [Fwd: Shift report 20031227 00:00-08:00]

From: Hironori Ito <hito@rcf.rhic.bnl.gov>
Date: Sat Dec 27 2003 - 07:38:06 EST

attached mail follows:


------------------------------------------------------------
Date: Sat Dec 27 07:33:14 2003, Shift: 00:00-08:00
Supervisor: Oana Ristea, with: Catalin Ristea
Log book#15 pages 17 - 17

Shift Summary:
------------------------------------------------------------
1. Beam from 1,30 to 5,30 when, just after gap cleaning, there was a QLI, which continues as we write.
2. run 9253 - D1=1/4B, D2=1/4B, D5=350B - 400k events.
   run 9254 - D1=1/4B, D2=1/4B, D5=350B - 325k events.
   run 9255 - pedestal for all detectors
   run 9256 - pedestal for TPCs.
3. No new problems found with detectors.

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Dec 27 07:38:39 2003

This archive was generated by hypermail 2.1.8 : Sat Dec 27 2003 - 07:38:57 EST